OpenStudy (anonymous):

What is the solution set of (3x - 1)2 = 5?

OpenStudy (anonymous):

i take it this is \[(3x-1)^2=5\] right?

OpenStudy (anonymous):

so \[3x-1=\pm\sqrt5\] and so you get \[3x=1\pm\sqrt5\] making \[x=\frac{1\pm\sqrt5}{2}\]

OpenStudy (anonymous):

oops typo there \[x=\frac{1\pm\sqrt5}{3}\]

OpenStudy (anonymous):

if you have \[(ax+b)^2=c\] do not multiply out and use the quadratic formula, that is working backwards